Describe what procedure you would follow to prepare a buffer solution to maintain a pH of 4.5?
1
Expert's answer
2017-05-08T08:27:09-0400
pH=pKk+lgVc/Vk

pKk=pKd(CH₃COOH)=4.756

Vc/Vk=x

4.5=4.756lgx

X=0.555

Vc/Vk=0/555

Vc=0.555Vk

100-Vk=0.555Vk

Vk=64.31

Vc=100-64.31=35.69